# Inferring Thunderstorm Occurrence from Vertical Profiles of Convection-Permitting Simulations: Physical Insights from a Physical Deep Learning Model

### v1

@@ 3.2 Sparse connections
In this section, we provide technical details on the sparse layer implementation. The input layer is given by an array of shape with a field dimension (iterating over the field profiles) and a height dimension (iterating over the vertical levels). Now, we consider a block of shape , where is the size of the block along the height dimension. We densely connect the nodes within this block to nodes in the following layer. Next, we slide the block by nodes along the height dimension and, again, densely connect the corresponding nodes to subsequent nodes of the following layer. This is schematically shown in Fig. 3. Starting with a block at the bottom of the input layer, we repeat this procedure until reaching the top of the input layer. Provided that is divisible by , this procedure leads to blocks and produces nodes in the following layer. We incorporate a shuffling mechanism that randomly permutes the order of the blocks for each example during training.
In contrast to a convolutional layer with a sliding kernel, all blocks in our sparse layer have their own set of free parameters. In total, the sparse layer contributes to the model. We have studied a large variety of -combinations and found that skill depends barely on the particular sliding block configuration as long as a sufficiently large number of parameters is exceeded. Setting , which corresponds to the smallest model configuration with saturating skill, we obtain blocks, and trainable parameters. In comparison to a fully dense layer, the parameter size is reduced by around .
@@ 3.3 Baseline model
In order to study the potential benefit of post-processing vertical NWP profiles instead of derived single-level predictors, we implement the ML model from Vahid Yousefnia et al. (2024), which we refer to as SALAMA 0D in the following (in contrast to SALAMA 1D, which processes one-dimensional vertical profiles). SALAMA 0D uses a set of 21 established predictors related to thunderstorm activity to infer calibrated probabilities of thunderstorm occurrence. The predictors are mapped to a single node via three dense layers with 20 nodes each. In fact, the architectures of SALAMA 0D and SALAMA 1D are identical in the last five layers (Fig. 2). Both models operate point by point, which allows for a direct and fair comparison. We retrain SALAMA 0D with training data from 2021 which matches the NWP model, the study region, and study period, used for SALAMA 1D. In contrast to the original paper, we feed the NWP data to SALAMA 0D on its native horizontal grid, reducing potential interpolation errors.
@@ 4.1 Model comparison
In order to get a first idea of the skill of the three models, we consider two cases with thunderstorm activity in Central Europe, namely July 24, 2023, 2200 UTC (case A), and August 2, 2023, 1200 UTC (case B). These two cases were chosen since they display multiple simultaneous convective regions of varying size. In Fig. 4, we show maps of the probability of thunderstorm occurrence for Central Europe as produced by the three models and compare them with lightning observations. The probability maps have been computed by retrieving the latest NWP forecast for each target time (case A: the 2-hour forecast of the 1500 UTC model run, case B: 0-hour forecast of the 1200 UTC run) and applying the SALAMA models to them. The ML models can be applied to all NWP ensemble members, producing separate probability output for each member. We show the results for only a single member.
@@
To compare the models quantitatively, we use the test set from Section 2. The first instrument we use for comparing model skill is a reliability diagram (Wilks, 2011; Bröcker and Smith, 2007). Partitioning the range of possible model probabilities into equidistant bins, we distribute the test examples among the bins according to the model probability they have been assigned. For each bin , we extract the observed relative frequency of thunderstorm occurrence, the bin-averaged model probability , and the number of examples per bin. A reliability diagram then consists of a calibration function and a refinement distribution. The calibration function is a plot of against , and measures whether the model probabilities are consistent with the observed relative frequency of thunderstorm occurrence, a characteristic known as reliability. A well-calibrated model exhibits a calibration function close to the 1:1 diagonal. The refinement distribution corresponds to the distribution of model probabilities. Skillful models are capable of producing well-calibrated model probabilities larger than climatology, which is referred to as resolution.
@@
In the upper panels of Fig. 5, we show for each of our models the corresponding reliability diagram with bins. All models display a similar degree of high reliability. We reiterate here that it is important to apply the analytic model calibration (2), otherwise high reliability could not be expected. The refinement distributions, as well, look similar. However, the model resolutions differ significantly: Following Vahid Yousefnia et al. (2024), we introduce the bin-wise contributions to reliability and resolution
@@
where denotes bin width. The area between and as a function of serves as a positively-oriented measure of skill. According to Murphy (1973), this area is equivalent to the Brier skill score (BSS). The lower panels of Fig. 5 display bin-wise reliability and resolution, revealing that the increase in skill, measured by the BSS, for the two SALAMA 1D models is attributed to improved resolution. For SALAMA 1D-2021, both low- and high-probability examples enhance resolution, while for SALAMA 1D-2022, all bins with contribute additional improvements to resolution.
@@
In Table 2, we summarize the performance of the three models using several additional skill scores established for problems with class imbalance (Wilks, 2011; Saito and Rehmsmeier, 2015; Vahid Yousefnia et al., 2024). These scores are positively oriented and bounded by unity. Across all skill scores, the SALAMA 1D models consistently outperform SALAMA 0D, with SALAMA 1D-2022 showing higher skill than SALAMA 1D-2021. It is also noteworthy that the inference time is comparable for all models; thus, the SALAMA 1D models are equally well-suited for real-time operational use, in addition to offering superior skill compared to SALAMA 0D.
| Skill score | 0D | 1D-2021 | 1D-2022 |
| Brier skill score (BSS) | |||
| Area under the precision–recall curve (PR-AUC) | |||
| Recall | |||
| Precision | |||
| -score | |||
| Critical success index (CSI) | |||
@@
So far, we have worked with a test set that consists of examples from NWP forecasts with a lead time of at most (Section 2). Next, we examine systematically how model skill depends on NWP forecast lead time. For this purpose, we generate test sets in which the examples result from NWP forecasts with a fixed lead time. As before, the position within the study region, the ensemble member index, and hour of the day, are randomly drawn from uniform distributions. The target times are drawn from the same test days as in Section 2.
@@
The lead-time dependence of skill is shown in Fig. 6. While we measure skill in terms of the BSS, we have checked that the results of this section do not qualitatively change when considering a different skill score. All three models exhibit an approximately exponential decrease in skill. The rate at which skill decreases is very similar for the three models. This suggests that the decrease of skill is not model-specific but results from an increasing NWP forecast uncertainty, which is consistent with previous work (Vahid Yousefnia et al., 2024). As a consequence, the SALAMA 1D models’ superior skill for low lead times is passed on to longer lead times.
@@
It is worth noting that the decrease in skill of SALAMA 0D is stronger than reported in Vahid Yousefnia et al. (2024). There, initial skill decreased by at most after , while the decrease here is approximately twice as much. This may result from using more diverse test sets in this study (we use twice as many test days to compile the training set). Furthermore, we do not apply any horizontal interpolation for compiling data sets in the present work, possibly allowing the model to rely on gridscale structures that would otherwise be masked by interpolation errors.
4.2 Sensitivity analysis
For the remainder of this section, we focus on SALAMA 1D-2022, referring to it simply as SALAMA 1D. In order to gain insight into how, given an input sample , the conditional probability of thunderstorm occurrence (SALAMA 1D model output) is inferred, we conduct a linear sensitivity analysis. For this purpose, we consider for a given scaled (Section 3) input sample the partial derivatives of :
@@ 4.2 Sensitivity analysis
Saliency varies from example to example and is used to interpret individual predictions. In order to gain more general insight, we propose considering statistical averages over suitable sets of examples. In the following, we evaluate saliency for the top and bottom probability percentile of test set examples. The resulting saliency maps characterize how SALAMA 1D tends to infer high and low probabilities of thunderstorm occurrence, respectively. We present them in Fig. 7 and Fig. 8, together with the corresponding average field profiles. Note that average saliencies of the different fields are stacked on top of each other. As a consequence, the saliency envelope quantifies how much individual height levels affect the model outcome. Average field profiles are shown in the same color as the corresponding average saliency. We left out the average pressure profile as it displays for both percentiles the same monotonically decreasing dependence of height, suggesting that pressure is essentially hydrostatic.
Beginning with the top percentile, we find one dominating peak at a height of , and a smaller one at . The large peak is almost entirely made up of contributions from temperature and pressure. To understand this, note that the model is not explicitly informed about the height of individual vertical levels; rather, these levels are randomly shuffled (Section 3). Due to the shuffling, we expect the pressure profile saliency to become a proxy for how much the model relies on vertical gradients. This is because hydrostatic pressure monotonically decreases with height and, therefore, contains information about vertical level spacings. Hence, the large peak suggests that the model reconstructs temperature lapse rates at the peak level. Additionally, Fig. 7 shows the average level of free convection (LFC) and equilibrium level (EL) for a mixed-layer parcel. The large peak is bounded by the LFC and the EL, which suggests that mixed-layer parcels are buoyant for heights in the vicinity of this peak. Positive buoyancy occurring in a conditionally unstable troposphere is known to be a crucial ingredient for thunderstorm development (Doswell et al., 1996) and is the basis for several traditional thunderstorm predictors, e.g. convective available potential energy. Conditional instability significantly depends on the environmental temperature lapse rate, which is consistent with temperature and pressure displaying high saliency between the LFC and the EL.
The highest-altitude peak closely coincides with the tropopause, which we see by computing the average tropopause height according to the WMO definition (World Meteorological Organization, 1957). While this peak receives similar contributions from different fields, the saliency of the cloud ice mixing ratio (QI) is maximal at this level. QI is present only at this height, as seen from the profile of average QI. This shows that the model actively takes ice particle content into consideration to infer thunderstorm occurrence. On the other hand, the model does not consider graupel content (QG) equally much. Presumably, this is because QG is maximal between the LFC and the EL, where the model has identified with the temperature lapse rate a more effective proxy for thunderstorm activity. As for specific humidity (QV), our study shows that the saliency of QV is highest around the LFC. Comparing the dew point profiles shows that high-probability examples exhibit more low-level moisture than low-probability examples. This supports the importance of low-level moisture for thunderstorm development, as dry air mixing in with rising air parcels reduces their buoyancy (Doswell et al., 1996; Markowski and Richardson, 2011).
Next, we turn to the bottom of examples. Note that the model is far less sensitive to input variability for the bottom class than for the top class—average saliency is roughly five orders of magnitude smaller. Nevertheless, average saliency varies as a function of height and field, displaying three distinct peaks. The upper two peaks coincide with those of the top class, though the contributions to the middle peak slightly change: The water quantities QC, QI, QG, as well as cloud cover (CLC) and vertical velocity (W) contribute more to this peak than for the top class. The profiles of these variables vanish in the vicinity of the middle peak for the bottom class but not for the top class, which suggests that the model infers thunderstorm occurrence from the absence of these fields.
The bottom class features an additional low-level peak at , which receives most contribution from temperature and pressure. As discussed before, this indicates that the model reconstructs a temperature lapse rate there, as well. Stable stratification in the lower troposphere is a known obstacle that parcels in the boundary layer need to overcome for thunderstorm development (Doswell et al., 1996). The energy required to overcome the barrier in parcel theory is given by convective inhibition. It depends on temperature, pressure, and specific humidity, which aligns with the composition of the peak.
A comparison between the CLC profiles of the two classes shows that the model associates non-vanishing cloud cover with a high probability of thunderstorm occurrence. In particular, close to the EL, CLC tends to be . This is consistent with anvil cloud top levels (Markowski and Richardson, 2011). Furthermore, it is noteworthy that high-probability examples tend to have northeastward wind profiles, whereas examples from the bottom class display southeastward winds. This is consistent with studies on the typical propagation direction of thunderstorms in Central Europe (Hagen and Finke, 1999).
The dominance of peaks related to environmental lapse rates and low-level moisture suggests that SALAMA 1D identifies thunderstorm occurrence essentially by patterns characteristic of convection initiation. These are typically mesoscale (Markowski and Richardson, 2011). Only to a lesser extent is model outcome affected by fine-grained (sub-mesoscale) patterns representative of ongoing convection, such as vertical velocity and liquid/ice particle content. This might seem surprising as we use simulation data from an NWP model capable of producing realistic-looking convection. However, predicting the correct location and timing of convection remains challenging for convection-permitting NWP models (Roberts and Lean, 2008). Therefore, SALAMA 1D may have learned to be cautious about fine-grained patterns if they do not reliably match lightning observations. As a result, it prefers larger-scale patterns, even though these broader structures can lead to more false alarms.

### v2

@@ 3.2 Sparse connections
In this section, we provide technical details on the sparse layer implementation, an illustration of which is provided by Fig. 3. The input layer is given by an array of shape with a field dimension (iterating over the field profiles) and a height dimension (iterating over the vertical levels). Now, we consider a block of shape , where is the size of the block along the height dimension. We densely connect the nodes within this block to nodes in the following layer. Next, we slide the block by nodes along the height dimension and, again, densely connect the corresponding nodes to subsequent nodes of the following layer. Starting with a block at the bottom of the input layer, we repeat this procedure until reaching the top of the input layer. Provided that is divisible by , this procedure leads to blocks and produces nodes in the following layer. We incorporate a shuffling mechanism that randomly permutes the order of the blocks for each example during training.
In contrast to a convolutional layer with a sliding kernel, all blocks in our sparse layer have their own set of free parameters. In total, the sparse layer contributes parameters to the model. We have studied a large variety of -combinations and found that skill depends barely on the particular sliding block configuration as long as a sufficiently large number of parameters is exceeded. Setting , which corresponds to the smallest model configuration with saturating skill, we obtain blocks, and trainable parameters. In comparison to a fully dense layer, the parameter size is reduced by around .
@@ 3.3 Baseline model
In order to study the potential benefit of post-processing vertical NWP profiles instead of derived single-level predictors, we implement the ML model from Vahid Yousefnia et al. (2024), which we refer to as SALAMA 0D in the following (in contrast to SALAMA 1D, which processes one-dimensional vertical profiles). SALAMA 0D uses a set of 21 established predictors related to thunderstorm activity to infer calibrated probabilities of thunderstorm occurrence. The predictors are mapped to a single node via three dense layers with 20 nodes each. The architecture of SALAMA 0D is the result of a hyperparameter study with increasingly many layers and nodes per layer until we observed saturation in terms of validation loss (Vahid Yousefnia et al., 2024). Therefore, SALAMA 0D skill is not limited by model architecture, and any difference in skill between SALAMA 1D and SALAMA 0D can be attributed to input data differences. Furthermore, both models operate point by point, which allows for a direct and fair comparison. We train SALAMA 0D with training data from 2021 which matches the NWP model, the study region, and study period, used for SALAMA 1D (Section 2). In contrast to the original paper, we feed the NWP data to SALAMA 0D on its native horizontal grid, reducing potential interpolation errors.
@@ 4.1 Model comparison
In order to get a first idea of the skill of the three models, we consider two cases with thunderstorm activity in Central Europe, namely July 24, 2023, 1700 UTC (case A), and August 2, 2023, 1200 UTC (case B). These two cases were chosen since they display multiple simultaneous convective regions of varying size. In Fig. 4, we show maps of the probability of thunderstorm occurrence for Central Europe as produced by the three models and compare them with lightning observations. The probability maps have been computed by retrieving the latest NWP forecast for each target time (case A: the 2-hour forecast of the 1500 UTC model run, case B: 0-hour forecast of the 1200 UTC run) and applying the SALAMA models to them. The ML models can be applied to all NWP ensemble members individually, producing separate probability output for each member. We show the results for only a single member. For both cases, we also show raw NWP output to see where the NWP model likely produces convection. To this end, we consider the column-maximal radar reflectivity product of ICON-D2-EPS. Specifically, for a given pixel, we compute the fraction of pixels within a radius of which exceed a threshold of (e.g., Theis et al., 2005; Roberts and Lean, 2008). Exceedance probabilities of reflectivity with thresholds between and have also been used in previous studies to identify thunderstorm occurrence (e.g., Mueller et al., 2003; Leinonen et al., 2022; Ortland et al., 2023).
@@
Raw NWP output structures have a smaller horizontal extension than the lightning regions in both cases, resulting in a considerable number of missed pixels even when the NWP model produces convection within a lightning region. The ML models tend to correct the size of the convective structures, with SALAMA 1D-2022 producing the least false alarms. Remarkably, the SALAMA 1D models can also produce high-probability output when lightning occurred but no convection has been triggered in the NWP model, as can be seen for the lightning regions over France for case A, which suggests that our ML models, SALAMA 1D-2022 in particular, may be able to correct for NWP model biases.
@@
To compare the models quantitatively, we use the test set from Section 22.3. The first instrument which we use for comparing model skill is a reliability diagram (Wilks, 2011; Bröcker and Smith, 2007). Partitioning the range of possible model probabilities into equidistant bins, we distribute the test examples among the bins according to the model probability they have been assigned. For each bin , we extract the observed relative frequency of thunderstorm occurrence, the bin-averaged model probability , and the number of examples per bin. A reliability diagram then consists of a calibration function and a refinement distribution. The calibration function is a plot of against , and measures whether the model probabilities are consistent with the observed relative frequency of thunderstorm occurrence, a characteristic known as reliability. A well-calibrated model exhibits a calibration function close to the 1:1 diagonal. The refinement distribution corresponds to the distribution of model probabilities. Skillful models are capable of producing well-calibrated model probabilities larger than climatology, which is referred to as resolution.
@@
In the upper panels of Fig. 5, we show for each of our models the corresponding reliability diagram with bins. All models display a similar degree of high reliability. We reiterate here that it is important to apply the analytic model calibration (2), otherwise high reliability could not be expected. The refinement distributions, as well, look similar. However, the model resolutions differ significantly: Following Vahid Yousefnia et al. (2024), we introduce the bin-wise contributions to resolution and reliability
@@
where denotes bin width. The bin-wise contributions to resolution are positively-oriented (“the higher, the better”), while the contributions to reliability are negatively-oriented. In consequence, the area between and as a function of serves as a positively-oriented measure of skill. According to Murphy (1973), this area is equivalent to the Brier skill score (BSS) if a random model based on climatology (Section 22.3) is applied as a reference. The lower panels of Fig. 5 display bin-wise reliability and resolution, revealing that the increase in skill, measured by the BSS, for the two SALAMA 1D models is attributed to improved resolution. For SALAMA 1D-2021, both low- and high-probability examples enhance resolution, while for SALAMA 1D-2022, all bins with contribute additional improvements to resolution.
@@
In Table 3, we summarize the performance of the three models using several additional skill scores established for problems with class imbalance (Wilks, 2011; Saito and Rehmsmeier, 2015; Vahid Yousefnia et al., 2024). These scores are positively oriented and bounded by unity. Across all skill scores, the SALAMA 1D models consistently outperform SALAMA 0D, with SALAMA 1D-2022 showing higher skill than SALAMA 1D-2021.
Skill score | 0D | 1D-21 | 1D-22 | 1D-21, 0D | 1D-22, 1D-21 | 1D-22, 0D |
BSS with climatology as reference | ||||||
Area under the precision–recall curve (PR-AUC) | ||||||
Recall | ||||||
Critical success index (CSI) | ||||||
@@
So far, we have worked with a test set that consists of examples from NWP forecasts with a lead time of at most (Section 22.1). Next, we examine systematically how model skill depends on NWP forecast lead time. For this purpose, we generate test sets in which the examples result from NWP forecasts with a fixed lead time. Just like before, these test sets with fixed lead times are sampled in a climatologically consistent manner and consist of examples. The target times are drawn from the same test days as in Section 22.3.
@@
The lead-time dependence of skill is shown in the upper panel of Fig. 6. While we measure skill in terms of the BSS, we have checked that the results of this section do not qualitatively change when considering a different skill score. All three models exhibit an approximately exponential decrease in skill. The rate at which skill decreases is very similar for the three models. This suggests that the decrease of skill is not model-specific but results from an increasing NWP forecast uncertainty, which is consistent with previous work (Vahid Yousefnia et al., 2024). As a consequence, the SALAMA 1D models’ superior skill for low lead times is passed on to longer lead times. In the lower panel of Fig. 6, we show the difference in skill as a function of lead time for all model pairs. Again, we find that the SALAMA 1D models consistently outperform SALAMA 0D at a confidence level of , with SALAMA 1D-2022 showing higher skill than SALAMA 1D-2021.
@@
It is worth noting that the decrease in skill of SALAMA 0D is stronger than reported in Vahid Yousefnia et al. (2024). There, initial skill decreased by at most after , while the decrease here is approximately twice as much. This may result from using more diverse test sets in this study (we use twice as many test days to compile the training set).
4.2 Interpretability study
For the remainder of this section, we focus on SALAMA 1D-2022, referring to it simply as SALAMA 1D. Our goal is to gain insight into how our model classifies input. We start by inspecting how the vertical profiles of the SALAMA 1D input fields look like on average for test set samples to which our model assigns a particularly high or low probability of thunderstorm occurrence. Figure 7, shows the average vertical profiles for the top and bottom probability percentile of test set samples. Shaded bands denote the symmetric confidence interval. Note that we converted specific humidity to dew-point temperature for a more straightforward comparison with temperature. For better orientation within the panels, we also plot the average tropopause height, which we compute according to the WMO definition (lowest level where the lapse rate drops to K/km, with the average lapse rate within km above remaining K/km, World Meteorological Organization, 1957).
The first column of panels in Fig. 7 shows temperature and dew-point temperature for the two percentiles. The top percentile troposphere displays more moisture than the bottom percentile, in particular in 2- height. This is consistent with the documented importance of moisture for thunderstorm development, as the buoyancy of rising air parcels is otherwise reduced by dry-air entrainment (Zhang et al., 2003; Peters et al., 2023; Marquis et al., 2023). In the second column in Fig. 7, we show average profiles of mixing ratios of cloud water (QC), cloud ice (QI), and graupel (QG). This suggests that the model uses non-vanishing profiles of QI and QG to discriminate between the top and the bottom percentile. For high-probability samples, ice particle content peaks close to the tropopause at a height of 10-, consistent with measurements of vertical hydrometeor distributions (e.g., Vivekanandan et al., 1999; Hubbert et al., 2018). The third column in Fig. 7 displays the average profiles of cloud cover (CLC). In general, our model associates non-vanishing CLC with a high probability of thunderstorm occurrence. In particular, close to the tropopause, CLC tends to be . This is consistent with anvil cloud top levels (Markowski and Richardson, 2010). In the fourth column in Fig. 7, we show vertical profiles of the three wind components U, V, and W. It is noteworthy that high-probability samples tend to have southwesterly wind profiles, whereas samples from the bottom percentile display northwesterly winds. This is consistent with studies on the typical propagation direction of thunderstorms in Central Europe (Hagen and Finke, 1999). On the other hand, vertical profiles of W vanish for both percentiles. We presume that due to convection displacement errors in the NWP model, the updraft regions within simulated deep convection rarely matches observed lightning observations. Therefore, SALAMA 1D may have learned not to rely on W for inferring thunderstorm occurrence. The last column of Fig. 7 shows the average profiles of pressure. Both profiles appear to be essentially hydrostatic. Surface pressure tends to be lower for the top percentile than for the bottom percentile.
Comparing the average profiles of the SALAMA 1D input fields for the two percentiles is useful to get a first idea whether our model separates the thunderstorm class from the majority class in a physically-interpretable manner. On the other hand, this analysis does not inform about the relative importance of the individual atmospheric variables. Therefore, we conduct a linear sensitivity analysis of the conditional probability of thunderstorm occurrence (SALAMA 1D model output) with respect to the input. The general idea is to consider for a given input sample the partial derivatives of :
@@ 4.2 Interpretability study
In order for saliency values to be comparable across all indices , we need to scale the input fields appropriately. This scaling accounts for the fact that the fields have different units and vary differently from one sample to another. Consider an unscaled input field (e.g., pressure), which we take as a function of height above ground. We define the corresponding scaled fields as
where and encode a characteristic background climatology for . We define them as
where stands for the th percentile of the variable in the brackets, evaluated for the training set. We then compute saliency with respect to the scaled fields.
While saliency varies from one sample to another and can be used to interpret individual predictions, we propose averaging over the top probability percentile to obtain more robust insight on the input fields and height levels which contribute most to high-probability output. To this end, we compute , where the angle brackets denote averaging over the top percentile samples. We take the absolute value (absolute saliency), as we consider feature importance to be linked to the (sample-averaged) intensity of the ML model’s linear response, irrespective of whether this response is positive or negative. In what follows, we refer to simply as saliency, i.e., with sample-averaging and taking the absolute value being implicitly implied unless stated otherwise. The resulting saliency maps for the two percentiles are shown in Fig. 8. Note that average saliencies of the different fields are stacked on top of each other. As a consequence, the saliency envelope quantifies how much individual height levels affect the model outcome.
The saliency envelope displays two distinct peaks, at and , respectively. The upper-level peak receives the most contributions by horizontal wind speed. Indeed, the saliencies of U and V are maximal near the tropopause, where the average horizontal wind velocity difference between the top and the bottom percentile is greatest (Fig. 7). This suggests that the model relies to a considerable extent on the learned climatological propagation direction of thunderstorms. Conversely, W saliency is approximately one order of magnitude smaller than the saliencies of U and V. QI saliency is maximal at the top of the troposphere, contributing significantly to the upper-level peak, as well. QI is present only at this height (Fig. 7), which suggests that the model actively takes ice particle content into consideration to infer thunderstorm occurrence. The only other hydrometeor field with significant non-vanishing saliency is specific humidity (QV). QV saliency peaks near the LFC, coinciding with the vicinity of maximal tropospheric moisture (Fig. 7). Similarly, CLC saliency is low but non-vanishing at all heights with non-zero CLC (Fig. 7).
Next, we turn to the mid-level peak. Apart from horizontal wind velocity, the mid-level peak receives most contributions from temperature. In addition, pressure saliency is non-vanishing only at the mid-level peak and near the surface. To understand this, note that our model is not informed about the height of individual vertical levels; rather, these levels are randomly shuffled (Section 33.1). Thus, our model can reliably infer heights levels—and in particular level spacings—only from pressure, which monotonously decreases with height (Fig. 7). Therefore, we expect pressure saliency to be a proxy for how much the model relies on vertical gradients. Since pressure saliency contributes to the mid-level peak and temperature saliency is high, we hypothesize that our model reconstructs mid-level temperature lapse rates. This is supported by the fact that the mid-level peak is bounded by the LFC and the EL of parcels lifted from the mixed layer, meaning that such parcels are buoyant for height levels in the vicinity of the mid-level peak. Positive buoyancy occurring in a conditionally unstable troposphere is known to be a crucial ingredient for thunderstorm development (Doswell et al., 1996) and constitutes the basis for several traditional thunderstorm predictors, such as convective available potential energy (CAPE). To test whether SALAMA 1D considers mid-level lapse rates, we show in Fig. 9a the distribution of 500- lapse rates for the top and bottom probability percentile. Indeed, essentially all high-probability samples are associated with conditionally unstable mid-level lapse rates, whereas the bottom percentile distribution extends further into absolutely stable mid-level lapse rates. Correspondingly, the top percentile profiles exhibit significantly more CAPE than those in the bottom percentile. (Fig. 9c). On the other hand, the distributions of the two percentiles show a significant overlap, which implies that considering mid-level lapse rates alone is not sufficient to infer a high probability of thunderstorm occurrence. As pressure saliency is non-vanishing also near the surface, we show in Fig. 9b distributions of near-surface (10-) lapse rates. Indeed, the distributions differ for the two percentiles. In contrast to before, near-surface lapse rates for the top percentile tend to be lower, and mostly absolutely stable, whereas lapse rates for the bottom percentile are mostly conditionally unstable. This is consistent with finding higher convective inhibition (CIN) in the top percentile than in the bottom percentile (Fig. 9d). This illustrates how the presence of CIN allows for CAPE to build up, which in turn makes convection initiation more likely (Carlson et al., 1983; Tuckman et al., 2023).
In summary, SALAMA 1D appears to rely on two categories of patterns. One category consists of patterns related to
horizontal wind direction,
mid-level and near-surface temperature lapse rates, or
low-level moisture.
We refer to these patterns as mesoscale since the underlying fields, such as temperature and pressure, vary more slowly in the horizontal than the fine-grained hydrometeor variables do. With the exception of horizontal wind direction, these patterns tend to be more characteristic of convective environments, in which convection is likely to be initiated, than of ongoing convection. In consistence with this, the top percentile samples show higher values of CAPE and CIN than the bottom percentiles do. Conversely, we identify a category of sub-mesoscale patterns, which relate to
tropopausal ice particle content,
vertical wind speed, or
cloud cover.
The sub-mesoscale patterns indicate ongoing convection. It is noteworthy that in terms of saliency, our model seems to rely less on sub-mesoscale patterns than on mesoscale patterns. As predicting the correct location and timing of convection remains challenging even for convection-permitting models (Roberts and Lean, 2008), SALAMA 1D may have learned to be cautious about sub-mesoscale patterns if they do not reliably match lightning observations. As a result, our model—on average—places greater emphasis (as indicated by saliency) on mesoscale patterns, while sub-mesoscale patterns serve to refine its skill. In practice, this also implies that SALAMA 1D is not strictly reliant on convection being produced in the NWP model. Instead, even if only some of the ingredients for convection are present, our model assigns a likely moderate but still well-calibrated probability of thunderstorm occurrence. In consequence, we expect some skill at correcting NWP output, as already suggested by the case studies in Fig. 4.